OpenStudy (anonymous):

find the x-coordinates of all points on the graph of: f(x)=6sinx-sin^2x at which the tangent line is horizontal. I get to deriving the function above then Im lost!

OpenStudy (anonymous):

*f(x)=6sinx-sin^2(x)

OpenStudy (anonymous):

once you get the derivative, set it equal to zero and solve

OpenStudy (anonymous):

my answers are in terms of pi/2, pi/3, pi/4 with options for 2npi or just npi

OpenStudy (anonymous):

\[\frac{d}{dx}6\sin(x)-\sin^2(x)=6\cos(x)-2\sin(x)\cos(x)\] set \[6\cos(x)-2\sin(x)\cos(x)=0\] then \[2\cos(x)(3-2\sin(x))=0\] so \[\cos(x)=0\] or \[2\sin(x)=3\] \[\sin(x)=\frac{3}{2}\] and then solve for x

OpenStudy (anonymous):

oops i made a mistake in factoring i think

OpenStudy (anonymous):

\[6\cos(x)-2\sin(x)\cos(x)=0\] \[2\cos(x)(3-\sin(x))=0\]

OpenStudy (anonymous):

so \[\cos(x)=0\] or \[\sin(x)=3\] which is not possible

OpenStudy (anonymous):

therefore \[\cos(x)=0,x=\frac{\pi}{2}+n\pi, n\in \mathbb Z\]

OpenStudy (anonymous):

Thanks for walking me through it, i really appreciate it!